OpenStudy (warmfridge):

I need help with the law of Cosine, can someone give me an example problem? It just takes a long time to do and i get lost easily.

OpenStudy (fortytherapper):

Hmm, Could you please think of two random numbers and one random angle for me?

OpenStudy (warmfridge):

11, 20, and 50 degrees

OpenStudy (fortytherapper):

So Law of Cosines is: \[c^2 = a^2+b^2 - 2abcos(C)\] You gave: \[a = 11\]\[b = 20\]\[C = 50\] We're looking for little c (the side of): Let's first plug in values that we have

OpenStudy (warmfridge):

11\[c ^{2}=11^2+20^{2}-2(11)(20)\cos(50)\]

OpenStudy (warmfridge):

ignore the 11 lol

OpenStudy (fortytherapper):

That's perfect. Now all we have to do is evaluate: \[11^2 + 20^2 = ?\]

OpenStudy (warmfridge):

521

OpenStudy (fortytherapper):

\[c^2 = 521 - 2(11)(20)\cos(50)\] I would next do 2(11)(20)

OpenStudy (warmfridge):

440

OpenStudy (fortytherapper):

Now to find what cos(50) is

OpenStudy (warmfridge):

.642?

OpenStudy (warmfridge):

feels wrong

OpenStudy (fortytherapper):

No, you're right .643. It's a calculus explanation but yeah lol. \[c^2 = (521) - (441)(.643)\] Now we just do some basic math steps

OpenStudy (warmfridge):

237.437

OpenStudy (fortytherapper):

\[c^2 = 237.437\] So last step to finding c is to root both sides, making it: \[c = \sqrt{237.437}\]

OpenStudy (warmfridge):

15.4

OpenStudy (fortytherapper):

Yep! That's the answer to that made up problem xP

OpenStudy (warmfridge):

thanks man

OpenStudy (fortytherapper):

You're welcome! You can use the same steps if they ask for a or b or the angle C